OpenStudy (anonymous):

which of the following is the equation for x: ax^2 = bx? x = a x = b x = 0 x = -b/a x = b/a

OpenStudy (fibonaccichick666):

so I assume they mean, for which of the x values listed below is the given function true

OpenStudy (fibonaccichick666):

so, what do you think?

OpenStudy (anonymous):

i think so?

OpenStudy (fibonaccichick666):

what value of x will make their statement true for any a,b?

OpenStudy (anonymous):

im not sure??

OpenStudy (anonymous):

\[ax^2-bx=0,x \left( ax-b \right)=0\] either x=0,or ax-b=0 ax=b x=b/a

OpenStudy (anonymous):

just those three answers? I'm just making sure.

OpenStudy (anonymous):

x=0,b/a

OpenStudy (anonymous):

no x = b?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

only x=0 x=b/a